1. Romy can climb a coconut tree at the rate of 10ft per minute and return at 20ft per minute. If it took him 3 minutes to climb and return. What is the height of the coconut?
2. There are two planes that fly in two directions. One is flying for East direction with the rate of 300kph and other is for West direction with the rate of 360kph. In how many minutes if they are 660 km apart?
3. Two planes travel from opposite direction. Plane A travel 70kph from one direction and plane B travel from opposite direction at the rate of 90kph. If the distance is 640km. In how many hours the two planes intersect?
4. Two planes x and y are 180 miles apart and mving from opposite direction at a rate of 20 and 30 miles per hour respectively. How many miles is covered by the slower plane when they intersect?
5. A man has 2 hours at his disposal. How far can he ride at the rate of 20 km per hour and return by walking at the rate of 5 km per hour.
I need this for me to pass my math subject.I hope for your detailed solution to these problems. Your response is highly appreciated. Thanks in advance!

Answer by mananth(16946)

1. Romy can climb a coconut tree at the rate of 10ft per minute and return at 20ft per minute. If it took him 3 minutes to climb and return. What is the height of the coconut?
let height of the tree be x feet.
Time = distance /speed
time to climb = x/10 minutes
time to return x/20 minutes
x/20 +x/10 = 3
LCD = 20
(x+2x )/20 =3
multiply by 20
3x= 60
/3
x= 20 feet height of tree.
.....
2. There are two planes that fly in two directions. One is flying for East direction with the rate of 300kph and other is for West direction with the rate of 360kph. In how many minutes if they are 660 km apart?
Effective speed of both planes = 300+360 kph.= 660kph since they are east west direction.
distance = 660 km apart
time = distance /speed.
660 /660 = 1 hour = 60 minutes
.....

3. Two planes travel from opposite direction. Plane A travel 70kph from one direction and plane B travel from opposite direction at the rate of 90kph. If the distance is 640km. In how many hours the two planes intersect?
they are traveling towards each other in a straight line.
effective speed = 70+90 = 160 kph
distance = 640 km
time = 640/160= 4 hours they take to meet.
....
4. Two planes x and y are 180 miles apart and mving from opposite direction at a rate of 20 and 30 miles per hour respectively. How many miles is covered by the slower plane when they intersect?
..
time taken by both the planes to meet is the same.
let it be t hours
rate of one plane = 20 mph
rate of other = 30 mph
distance traveled by one is 20t and the the 30 t
20t+30t = 180
50 t= 180
/50
t=180/50
t=1.6 hours.
..
speed of slower plane = 20 mph.
time t = 1.6 hours
so distance traveled = 20*1.6 = 32 miles.

5. A man has 2 hours at his disposal. How far can he ride at the rate of 20 km per hour and return by walking at the rate of 5 km per hour.
let the distance he can go be x km
riding time will be x/20 hours
walking time will be x/5 hours
..
both these times put together = 2 hours
x/20 + x/5 = 2
LCD = 20
x+4x / 20 = 2
5x= 40
x= 8 km he can go 8 km.

wish you good luck
I need this for me to pass my math subject.I hope for your detailed solution to these problems. Your response is highly appreciated. Thanks in advance!